OpenStudy (hkmiu):

What are the first three terms of the arithmetic series if a1 = -1, an = -115, and Sn = -1160?

OpenStudy (anonymous):

isn't the formula for the sum something like \(S=\frac{n}{2}(a_1+a_n)\)?

OpenStudy (anonymous):

i think that is right

OpenStudy (anonymous):

so we can solve for \(n\) \[\frac{n}{2}(-1-115)=-1160\] or \[\frac{n}{2}\times (-116)=-1160\]

OpenStudy (anonymous):

divide by \(-116\) to get \(\frac{n}{2}=10\) and so \(n=20\)

OpenStudy (anonymous):

Are you sure satellite? I don't remember that...

OpenStudy (anonymous):

i could be wrong

OpenStudy (anonymous):

No that's right.

OpenStudy (anonymous):

I was thinking to find any term in an arithmetic sequence.

OpenStudy (anonymous):

i could be wrong, but i am not http://en.wikipedia.org/wiki/Arithmetic_progression

OpenStudy (anonymous):

in any case that does not answer the question, because the question is not "how many terms are there" but "what are the first 3 terms" so there is more work to do

OpenStudy (hkmiu):

Whoops i forgot to check on this.

OpenStudy (anonymous):

lol you are here now right? we are not done

OpenStudy (hkmiu):

LOL, I'm here. I'm reading and catching up!

OpenStudy (anonymous):

ok while you catch up i will see if i can figure out how, now that we know \(n=20\) to find \(d\)

OpenStudy (anonymous):

ah yes, \(a_n=a_1+(n-1)d\)

OpenStudy (anonymous):

let me know when you catch up

OpenStudy (anonymous):

yeah, that's the one :P .

OpenStudy (hkmiu):

I caught up. I was abot to write the formula thn you got it :P

OpenStudy (hkmiu):

sooo its 1 + ( 20 - 1)d = an

OpenStudy (anonymous):

ok so we can solve for \(d\) right?

OpenStudy (hkmiu):

Yesss! I can pick it up from here. Thank you! :)

OpenStudy (anonymous):

hold on i think you have a mistake there

OpenStudy (hkmiu):

I mean a1 = -1

OpenStudy (anonymous):

\[-1 + ( 20 - 1)d = -115\]

OpenStudy (anonymous):

oh ok good

OpenStudy (hkmiu):

Hahahaaa ^____^ thank youuu!

OpenStudy (anonymous):

yw